OpenStudy (anonymous):

what are the roots of this equation, please show all working. r^2 + r - 2 =0

OpenStudy (anonymous):

b^2 - 4ac = 1 - 4*1*(-2) = 9

OpenStudy (anonymous):

or factoring r^2 + r - 2 = (r + 2)*(r - 1)

OpenStudy (anonymous):

-> r = -2 or r = 1

OpenStudy (anonymous):

thankyou very much i was after the factoring answer